Assume that it takes 49.61 min to produce the 20th unit and 40.03 min to produce the 50thunit. What is the learning curve? How many cycles are needed to reach a standard time of 35?
Answer #1

The formula that will be used here is: Tn = T1*C

where Tn = time required for nth unit, T1 = time required for 1st unit and C = learning curve coefficient

For the 20th unit: T20 = T1*C for 20 or 49.61 = T1*C for 20. Thus T1 = 49.61/C for 20

For 50th unit: T50 = T1*C for 50 or 40.03 = T1*C for 50. Thus T1 = 40.03/C for 50

Now 49.61/C for 20 = 40.03/C for 50

or 1.2393 = C for 20/C for 50

We know that C = N ^b where b is the log of the learning rate/log of 2

Thus 20^b/50^b = 1.2393

or b = -0.2342

also b = log of learning rate/log 2

Solving we get learning rate = 0.85

Thus learning curve = 85%

Let the no. of cycles be "x".

Thus 35 = T1*C

Now 49.61 = T1*C for 20 at 85%

or 49.61 = T1*0.495

or T1 = 49.61/0.495 = 100.22 minutes

So 35 = T1*C

or 35 = 100.22*C

or C = 35/100.22 = 0.349

C = N^b

Thus 0.349 = N^-0.2342

or N = 89

Thus standard time of 35 will be reached in the 89th unit.